OpenStudy (anonymous):

f(x)= -2x^2+2x+2 find the vertex and line of symetry

OpenStudy (anonymous):

Complete the square to put it in vertex form.

OpenStudy (anonymous):

i think the x coordinate would be 1/2 which would also be the line of symetry right?

OpenStudy (anonymous):

not sure what the y coordinate would be

OpenStudy (anonymous):

Plug in 1/2 for x, and you can find y.

OpenStudy (anonymous):

= -2(x^2 - x - 1) = -2 ((x- 1/2)^2 - 1/4 - 1) = -2(x- 1/2)^2 - 5/4) vertex is at (1/2, 2 1/2)

OpenStudy (anonymous):

line of symmetry is x = 1/2